OpenStudy (anonymous):

Solve the equation 2/3(x − 4) = 2x for x. 2 −2 −8 −4

OpenStudy (anonymous):

2/3x-8/3=2x

OpenStudy (anonymous):

3/2* [2x] = x-4 3x = x-4 2x=-4 x=-2

OpenStudy (anonymous):